OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (mathmate):

The set is {-1,1} An operation is closed when the operation is operated on \(any\) members of the set will give a result which is within the set. For example, for the addition operation, -1+-1 =-2 Not in the set, so the operation is not closed on the given set {-1,1}. Check with other operations to find the required answer.

OpenStudy (anonymous):

huh?

OpenStudy (michele_laino):

for example, your set is not closed with respect to addition, since: \(1+(-1)=0\) and \(0\) doesn't belong to your set

OpenStudy (anonymous):

ok, so addition?

OpenStudy (michele_laino):

so, addition is a wrong option

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

multiplucation

OpenStudy (michele_laino):

we have to check for multiplication: so, \(1 \cdot (-1)=(-1) \cdot 1=-1\) and \(-1\) belongs to your set \(1 \cdot 1=1\) and \(1\) belongs to your set \((-1) \cdot (-1)=1\) and \(1\) belongs to your set so, what can you conclude?

OpenStudy (anonymous):

multiplication is right

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

ok, would division also work

OpenStudy (michele_laino):

yes! also division works!

OpenStudy (anonymous):

yay, awesome thank you so much

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (michele_laino):

ok! Please open a new question

OpenStudy (anonymous):

ok